# arXiv:2602.23010v2[cs.GR] 01 Mar 2026

### v2

@@ 9 Conclusion
We have presented Helmlab, a 72-parameter analytical color space that achieves stress 23.30 on combvd, which is 20.2% lower than CIEDE2000. The achromatic axis is guaranteed to have $C<10^{-6}$ via the neutral correction, and a rigid rotation brings hue error down to 18.1° RMS without affecting the distance metric. A blue-band refit with sub-dataset penalties reduces gradient non-uniformity in the blue–cyan region by $8.9\times$ at a cost of only +0.08+0.08 stress points, while improving stress on 5 of 6 sub-datasets. The neutral correction is the key element that makes the space usable for both measurement and generation: it lets the optimizer find its best parameters freely and then fixes the achromatic axis. Code, trained parameters, and utilities are available at https://github.com/Grkmyldz148/helmlab. An interactive demo is at https://grkmyldz148.github.io/helmlab/demo.html and documentation at https://grkmyldz148.github.io/helmlab/.
@@ Acknowledgments
Thanks to the 71 anonymous observers who participated in the bidirectional evaluation study during development.
Appendix A Parameter Table
Table˜5 lists all 72 optimized parameters grouped by pipeline stage.
Table 5: Optimized Helmlab parameters (v20b, $\varphi=-28.2$°).
@@ Appendix A Parameter Table
$\mathbf{M}_{1}$ (9) Row 0 [0.734, 0.240, -0.158] XYZ $\to$ cone-like
Row 1 [-0.330, 1.235, -0.000]
Row 2 [0.086, 0.341, 0.832]
$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.389, 0.416, 0.424 Power compression
$\mathbf{M}_{2}$ (9) Row 0 [-0.420, 0.485, 0.642] Compressed $\to$ Lab
Row 1 [1.933, -2.718, 0.764]
Row 2 [0.006, 1.628, -1.261]
Hue corr. (8) $\alpha_{1..4},\beta_{1..4}$ see JSON 4-harmonic Fourier
H-K (6) $w,p,m,s_{1},c_{2},s_{2}$ 0.389, 0.849, -0.494, … Chroma $\to$ L
Light. (8) $p_{1},p_{2},p_{3},Lh_{c1},Lh_{s1},\lambda_{d},h_{c},h_{s}$ 0.222, -0.727, 0.695, … Cubic + Lh + dark
Chroma (18) 4-harm + power + $lc_{1},lc_{2}$ + HLC see JSON Hue/power/L-dep
Hue-L (4) $g_{c1..2},g_{s1..2}$ see JSON Hue $\to$ L
Distance (7) $s_{L}$ $1.01\times 10^{-3}$ Pair-dep L weight
$s_{C}$ 0.022 Pair-dep C weight
pp 0.804 Minkowski exponent
$w_{C}$ 1.046 Chroma weight
$c$ 1.590 Compression
$q$ 1.100 Post-power
$\alpha$ 0.000 Linear term
Full parameter values are available in the accompanying metric_params.json file.

### v3

@@ 10 Conclusion
[…]
The neutral correction is the architectural element that lets the two specialisations co-exist: MetricSpace can chase the distance-fit basin freely, because the achromatic axis is then restored exactly by post-correction; GenSpace achieves the same guarantee structurally, by construction. Together with the rigid rotation invariance, this lets us decouple “what is the best distance metric” from “what should the chroma plane look like for generation”.
Helmlab is shipped in production: helmlab on PyPI and npm, four spaces and a $\Delta E$ method merged into Color.js (PR #722), and a PostCSS plugin (postcss-helmlab) for build-time CSS function transformation. Source, trained parameters, and a full ColorBench-driven benchmark report are at https://github.com/Grkmyldz148/helmlab. An interactive demo is at https://grkmyldz148.github.io/helmlab/demo.html and documentation at https://grkmyldz148.github.io/helmlab/.
@@ Acknowledgments
We thank the 71 observers — listed by first name or chosen handle in the released dataset (datasets/human_feedback.json, github.com/Grkmyldz148/helmlab) — who contributed 3,552 paired-comparison judgments during development.
We also gratefully acknowledge the open-source review of the underlying PR #722 on color-js/color.js: in particular @facelessuser (near-black achromatic instability, M2 normalisation regressions, the clean 64-bit matrix derivation methodology), @lloydk (the WebGPU chart library that surfaced the h $\approx$ 264° blue fold and related cubic-cusp artefacts), @Userminusone (the Metal-shader hue/ lightness probe, scale-invariance critique of NR, the boundary-warp straightening alternative), and @svgeesus (the original paper proposal, the “0% should be achromatic” authoring convention, and the ICC Displays working-group framing). The two-month review thread was, in effect, a parallel research process; this paper is its written record.
Appendix A Parameter Table (MetricSpace v21)
Table˜6 lists all 72 trained MetricSpace v21 parameters grouped by pipeline stage. The architecturally-reserved 22 surround/hue-modulated parameters are held at zero in v21 and omitted from this table; their full list is in the JSON checkpoint.
Table 6: MetricSpace v21 parameters ($\varphi=-28.2$°). Trained on combvd + He 2022 + MacAdam 1974 with sub-dataset-balanced loss (Equation˜16).
@@ Appendix A Parameter Table (MetricSpace v21)
$\mathbf{M}_{1}$ (9) Row 0 [-0.7213, -0.4534, -0.1929] XYZ $\to$ cone-like
Row 1 [-0.7882, -1.7952, -0.0876]
Row 2 [-0.0918, -0.4577, -1.2922]
$\gamma$ (3) $\gamma_{0},\gamma_{1},\gamma_{2}$ 0.4723, 0.5149, 0.5113 Signed power compression
$\mathbf{M}_{2}$ (9) Row 0 [-0.2636, -0.4168, -0.4927] Compressed $\to$ Lab
Row 1 [-1.8898, -3.1212, -1.0422]
Row 2 [-0.3585, -1.7694, -1.4121]
Hue corr. (8) cos1..4, sin1..4 see JSON 4-harmonic Fourier rotation $\delta(h)$
H-K (6) $w_{\mathrm{HK}},p_{\mathrm{HK}},\mathrm{hue\_mod}$ 0.2676, 0.8935, 0.7173 Base + envelope
sin1, cos2, sin2 0.6915, 0.4865, 0.9853 2-harmonic hue mod $f_{\mathrm{HK}}(h)$
Lightness (8) $p_{1},p_{2},p_{3}$ 0.5385, 0.1251, 0.6769 Cubic correction
$Lh_{c1},Lh_{s1}$ -0.4963, -0.0956 Hue-modulated additive
$\lambda_{d},h_{c},h_{s}$ -0.0291, 1.3347, -0.1699 Dark-region exp compression
Chroma (18) cs1..4, cp1..2, $lc_{1,2}$, hlc1,2 see JSON 4-harm CS + 2-harm CP + L-dep + HLC
Hue-L (4) $g_{c1..2},g_{s1..2}$ see JSON 2-harmonic $L\mathrel{*}\!\!=\!\exp(g(h))$
Distance (7) $s_{L}$ -0.9155 Pair-dep L weight (Equation˜12)
$s_{C}$ -2.9268 Pair-dep C weight (Equation˜13)
$p$ -1.9737 Minkowski exponent
$w_{C}$ -3.9660 Chroma weight
$c$ -52.473 Compression
$q$ -0.4790 Post-compress power
$\alpha$ -0.0000 Linear-asymptote term (unused in v21)
Full parameter values, including the architecturally-reserved zero slots and all chroma-stage Fourier coefficients, are available in research/checkpoints/metricspace_v21.json.
Appendix B Parameter Table (GenSpace v0.11.1)
Table 7: GenSpace v0.11.1 parameters.
Stage Value
$\mathbf{M}_{1}$ $\times3$, see JSON (Bradford CAT pre-baked)
Transfer depcubic, $\alpha=0.021$
$\mathbf{M}_{2}$ $\times3$, see JSON
Chroma power 0.978
PW $L$ correction 19-point dense, see JSON
Enrichment L-gated hue rotation
amplitude 0.058
center hue 264.5°
$\sigma$ 0.7 (rad)
gate $L$ window [0.37, 1.0]
Full parameter values ship bundled with the helmlab package (helmlab.GenSpace().params) and are written out as src/helmlab/data/gen_params.json in the source distribution. White(D65) maps to $L=1$ exactly, the achromatic axis is held at $|a|,|b|<10^{-4}$, and the structurally-achromatic property holds because the depcubic transfer is signed-symmetric and the smooth neutral-blend in Stage 2 collapses sub-microscopic channel spread to the channel mean before $\mathbf{M}_{2}$ projection.